# Continuity and equivariant dimension

### v2

@@ 3 Local-Triviality Dimension and Fields
and satisfy the conditions that each $\alpha_{i,s}^{k}$ is a projection and $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{i,s}^{*}\alpha_{i,s}$ is invertible on each $U_{i}$. As before, we have that the $\alpha_{i,s}$ are in the generating spectral subspace of the ${\mathbb{Z}}/k$ action by design. However, each $\alpha_{i,s}$ is only a section on $U_{i}$, not the full space $X$.
@@
A partition of unity argument (see [34, Theorem 36.1]) allows us to rescale the operators $\alpha_{i,s}\in\Gamma({\mathcal{M}}|_{U_{i}})$ into global sections $\phi_{i}^{1/2}\alpha_{i,s}\in\Gamma(M)$, where the $\phi_{i}$ are nonnegative scalar-valued continuous functions on $X$, supported in $U_{i}$, with $\phi_{i}|_{V_{i}}\equiv 1$ and $\sum\limits_{i=0}^{\operatorname{ix}}\phi_{i}\equiv 1$. Note that each $\phi_{i}$ is a central element of $\Gamma({\mathcal{M}})$ and is fixed by the action of ${\mathbb{Z}}/k$, so multiplying by $\phi_{i}^{1/2}$ has not removed any of the key properties of $\alpha_{i,s}$. Because $1\leq s\leq\operatorname{\mathrm{qt}}$ and $0\leq i\leq\operatorname{ix}$, we have produced a list of $\operatorname{\mathrm{qt}}\cdot(\operatorname{ix}+1)$ normal operators $\alpha_{i,s}$ in the generating spectral subspace of the ${\mathbb{Z}}/k$ action such that $\sum\limits_{i,s}\alpha_{i,s}^{*}\alpha_{i,s}$ is (globally) invertible. By Proposition 3.5, we have that
@@
The implication a $\Longrightarrow$ b follows from the upper bound 3-4 on $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))$, where we note that $\operatorname{ix}({\mathcal{M}},\triangleright)$ is always finite, and the full spectrum condition of (a) guarantees that the definition of $\operatorname{\mathrm{qt}}$ does not involve dividing by zero. Moreover, b $\Longrightarrow$ c follows from the general result [19, Theorem 3.8]. As for c $\Longrightarrow$ a, freeness means by [33, Theorem 8.1.7] that the ${\mathbb{Z}}/k$-grading on $A:=\Gamma({\mathcal{M}})$ induced by the action is strong:
@@
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A)=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A)=1\quad\text{and}\quad\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{SLT}}(A)=\infty$
@@
Example 3.7 already argues that none of the dimensions can vanish, and the upper bound of $1$ for the ordinary (hence also weak) LT dimension follows from Theorem 3.15a by the selfsame example: $\operatorname{ix}({\mathcal{M}},\triangleright)=1$, since a cover by two punctures of ${\mathbb{P}}^{1}$ trivializes everything in sight.
As for the infinitude of the strong LT dimension, recall first that the action at $x\in{\mathbb{P}}^{1}$ is conjugation by the self-adjoint (hence involutive) unitary having $x$ (regarded as a line in ${\mathbb{C}}^{2}$) and the orthogonal complement $x^{\perp}\leq{\mathbb{C}}^{2}$ as its $1$- and $(-1)$-eigenspaces, respectively.
@@
[…]
real projective plane $\cong\quad\text{real projective plane }{\mathbb{R}}{\mathbb{P}}^{2}.$
Because (by oddness) each eigenspace of $a_{i}$ is orthogonal to each eigenspace of $u$, 3-5 coincides with the canonical surjection ${\mathbb{S}}^{2}\to{\mathbb{R}}{\mathbb{P}}^{2}$ nowhere. A continuous lift ${\mathbb{S}}^{2}\to{\mathbb{S}}^{2}$ thereof (which does exist by [34, Lemma 79.1], as ${\mathbb{S}}^{2}$ is simply-connected) will thus have no coincidences with either the identity or the antipodal map, once more violating [58, Corollary 1.24]. $\blacksquare$
@@
by projectivity, as in the proof of [12, Proposition 5.4], via [45, Proposition 2.10]. Define $a:=\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t\otimes 1)$. Note that $\pi_{x}(a)$ is invertible, so we may choose an element $b\in A$ such that $\pi_{x}(b)=(\pi_{x}(a))^{-1}$. That is, both $\|ab-1\|_{x}=0$ and $\|ba-1\|_{x}=0$. By [47, Proposition 1.2], for each $c\in A$ the mapping
@@
Consider the antipodal ${\mathbb{Z}}/2$-action on the $\theta$-deformed spheres $C(\mathbb{S}^{3}_{\theta})$, which form a continuous field. The Borsuk-Ulam theorem implies that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(C({\mathbb{S}}^{3}))=3$. On the other hand, for $\theta\not\in{\mathbb{Z}}$ the proof of [39, Theorem 2.10] shows that if $C({\mathbb{S}}^{3}_{\theta})$ is noncommutative, with $z_{j}=x_{j}+iy_{j}$, then
@@
We conclude that the weak local triviality dimension function is upper semicontinuous, but not continuous, in the parameter $\theta$. $\blacklozenge$
@@
where ${\mathbb{S}}^{1}$ is a maximal torus: see [29, Theorems 1 and 2] and [28, Theorem III.1.11.1], for $G=SU(2)$, as recalled in [51, Introduction].
@@
    $SU(2)\xrightarrow{\quad}PSU(2)\cong SO(3);and$ (3-6)
@@
    resulting from realizing ${\mathbb{C}}^{n}$ as the irreducible $n$-dimensional $SU(2)$-representation (see [54, Theorem VIII.4.2]).
@@
On the other hand, $\sigma$ acts on ${\mathbb{C}}^{n}$ with eigenvalues ±$\pm\lambda$, with $\lambda=1$ or $\lambda=i$ depending on the parity of $n$, and with the (±$(\pm\lambda)$-eigenspaces of dimension as close to equal as possible. It follows that $\sigma$ induces the usual gauge $({\mathbb{Z}}/2)$-action on $M_{n}$ resulting from the latter’s realization as the graph algebra of a length-$n$ directed path. Finally, [12, Proposition 4.5] shows that
@@
The $A_{\theta}$ are the fibers of a continuous field by [50, Theorem 8.13 and Example 10.2], and all admit ergodic actions (see [20, p.537]) by the ordinary 2-torus ${\mathbb{T}}^{2}\cong({\mathbb{S}}^{1})^{2}$. This constitutes a continuous field of actions in the sense of [47, Definition 3.1]. Note the following.
@@
  • On the other hand, for non-integral values of $\theta$, none of the dimensions can vanish. The ${\mathbb{T}}^{2}$-action is ergodic, so the eigenspaces for the characters of ${\mathbb{T}}^{2}$ are all 1-dimensional (see [59, Theorem 1]). Therefore an equivariant morphism
@@
[…]
@@
with full equivalence when ${\mathbb{G}}$ is of Kac type [36, post Proposition 1.7.9].
@@
Recall from [60, Corollary 4.1] that $\varphi$ uniquely retrieves a completely positive contractive order-zero [60, Definition 2.3] map
@@
Note that $\widetilde{\varphi}$ is ${\mathbb{G}}$-equivariant in the usual sense, and by [60, Theorem 3.3] it is a scaling of a morphism $C({\mathbb{G}})\to A$ by the ${\mathbb{G}}$-invariant, positive, non-zero element $h:=\widetilde{\varphi}(1)\in A$. Since ergodicity of the action implies that $h$ is a scalar, we can divide by it to recover a ${\mathbb{G}}$-equivariant $C^{*}$-morphism $C({\mathbb{G}})\to A$.
Observe also that a ${\mathbb{G}}$-equivariant $C^{*}$-morphism is by necessity one-to-one on the dense Hopf *-algebra [36, §1.6] ${\mathcal{O}}({\mathbb{G}})\leq C({\mathbb{G}})$: the ${\mathbb{G}}$-invariant state
@@
restricts to $h_{{\mathbb{G}}}$ along the assumed morphism $C({\mathbb{G}})\to A$, and $h_{{\mathbb{G}}}$ is faithful on ${\mathcal{O}}({\mathbb{G}})$ [36, Corollary 1.7.5]. If the action is 0-dimensional, then, the multiplicity of a ${\mathbb{G}}$-representation $\rho$ in $A$ is precisely $\dim\rho$. For ergodic actions of Kac groups this is also an upper bound for that multiplicity [7, Theorem 17] (see also [59, Theorem 1] for classical compact groups), hence the implication a $\Rightarrow$ d for Kac-type ${\mathbb{G}}$. $\blacksquare$
@@
of ergodic ${\mathbb{T}}^{n}$-actions in Example 3.20 defined over the skew-symmetric matrices $\theta$ has local-triviality dimension equal to 00 precisely when $\theta$ is integral. Moreover, it has local-triviality dimension $\infty$ elsewhere. This proves a stronger claim than used in Example 3.20.
@@
  • if the weak local-triviality dimension (or rather, any one of the local-triviality dimensions) of an action is finite, then [19, Theorem 3.8] shows that the action is free;
@@
respectively for some primitive $q^{th}$ root of unity $\zeta$ (see [20, proof of Proposition 12.2, (12.7)]). This in fact realizes $M_{q}$ as a cocycle twist [17, §2.2] $C^{*}(\Gamma,\sigma)$ of the group algebra of
@@
by [52, Theorem 9.74 and Theorem following Example 9.65]. Next, [33, Theorem 8.1.7] once more shows that the induced grading by the Pontryagin-dual group $\Gamma$ is strong:
@@ 4 Rational tori and spheres
However, if some particular $t_{j}$ equals 00, then $z_{j}=t_{j}u_{j}=0$ is never changing, so the function $g$ must not depend on $u_{j}$. That is, $g$ is in the $C^{*}$-algebra generated by the other coordinate functions. This perspective extends to the $\theta$-spheres and quantum tori, as in [35, Theorem 2.5]. We recall this result and introduce some notation below.
@@
    for the “first-quadrant” portion of the $(n-1)$-sphere.
@@
    Each ${\mathbb{S}}^{n-1}_{++,F}$ is the interior of an $(|F|-1)$-dimensional simplex, and the closed simplex ${\mathbb{S}}^{n-1}_{+}$ is the disjoint union of the various open simplices ${\mathbb{S}}^{n-1}_{++,F}$.
@@
    ${\mathbb{S}}^{n-1}_{+,F}$ $:=\,\,\overline{{\mathbb{S}}^{n-1}_{++,F}}$
    =$=\,\,\left\{(t_{i})\in{\mathbb{S}}^{n-1}_{+}\ |\,\,\forall i\not\in F,\ t_{i}=0\,\right\}.$
For a skew-symmetric matrix $\theta\in M_{n}({\mathbb{R}})$, recall from [35, Theorem 2.5] that
@@
For $k=2n-1$, each generator $z_{j}=x_{j}+iy_{j}$ of $C({\mathbb{S}}^{2n-1}_{\theta})$ is normal and odd, so the identity $1=\sum\limits_{j=1}^{n}z_{j}z_{j}^{*}=\sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}$ shows that the weak and plain local-triviality dimensions are at most $12n-$ 1. A similar argument applies to $k=2n-2$, with the final generator $z_{n}=x_{n}$ self-adjoint. $\blacksquare$
@@
Every $\theta$-sphere has $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))\geq k$. If each entry $\theta_{jk}$ is a rational number with odd denominator, then in fact $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))=k$.
@@
The noncommutative Borsuk-Ulam theorem [39, Corollary 3.8] implies that if $m<k$, there is no $\mathbb{Z}/2$-equivariant morphism $C({\mathbb{S}}^{m}_{\phi})\to C({\mathbb{S}}^{k}_{\theta})$. This applies in particular to the case when the domain sphere is commutative. Nonexistence of an equivariant morphism $C({\mathbb{S}}^{m})\to C({\mathbb{S}}^{k}_{\theta})$ for $m<k$ is equivalent to the claim $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))\geq k$.
Now, assume that each $\theta_{jk}$ is a rational number with odd denominator, and suppose $k=2n-1$ is odd. There exists an odd positive integer $M$ such that in the associated quantum torus $C(\mathbb{T}^{n}_{\theta})$, the elements $u_{1}^{M},\ldots,u_{n}^{M}$ are central. The functions $f_{i}:{\mathbb{S}}^{n-1}_{+}\to C(\mathbb{T}^{n}_{\theta})$ defined by
@@
written in lowest terms. From [46, Theorem 3.12] (or [20, Proposition 12.2]) we see that, in this case, there is a rank-$q$ vector bundle ${\mathcal{E}}$ on ${\mathbb{T}}^{2}$ such that
@@
This is also recalled in [26, §2.5], which is perhaps more accessible. The center
@@
is generated by the two unitaries $u^{q}$ and $v^{q}$ [20, Corollary 12.3], and $A_{\theta}$ is a bundle of $q\times q$ matrix algebras over that center. This also makes $A_{\theta}$ into a rank-$q^{2}$ Azumaya algebra (see [32, §§13.7.6 and 13.7.13], [9, §§III.1.3 and III.1.4], [1, Definition 5.4.17], etc.): it is finitely-generated projective over its center $Z$, with
@@
    The discussion following [26, Exercise 2.7] claims that the rank-$q$ vector bundle ${\mathcal{E}}$ on ${\mathbb{T}}^{2}$ yielding 4-3 is flat, which it cannot be. A flat bundle (see [27, §1.2]) admits a connection with vanishing curvature, or local trivializations with locally constant transition functions, and flat bundles have vanishing Chern classes by [27, Proposition II.3.1]. In contrast, the Chern class
@@
    is identified in [46, Proposition 3.8 and Theorem 3.9] with an integer $a$ for which $ap+bq=1$ (the twist of [46, p.299] is $-a$). In particular, that Chern class cannot vanish. Note that the vector bundle ${\mathcal{E}}$ is only unique up to the tensor product with a line bundle [15, Théorème 9], so $a$ is only unique modulo $q$. This is also compatible with the fact that $A_{\theta}$ only depends on the coset of $\theta$ modulo ${\mathbb{Z}}$.
    Additionally, vector bundles over compact connected Riemann surfaces are completely determined by their ranks and degrees (i.e. Chern classes $c_{1}$) by [57, Proposition, p.2]. Flat bundles on ${\mathbb{T}}^{2}$, for that reason, are topologically trivial. They need not be holomorphically trivial, but ‘topologically’ is what matters in the present context.
@@
    What does hold, and what was perhaps meant by the flatness claim in the paragraph preceding [26, Exercise 2.8], is that the endomorphism bundle ${\mathcal{E}}\otimes{\mathcal{E}}^{*}$ (recovering $A_{\theta}$ as the section algebra 4-3) is flat, and in fact is topologically trivial as a vector bundle (but not as a bundle of algebras!). The first Chern class $c_{1}({\mathcal{E}}\otimes{\mathcal{E}}^{*})\in H^{2}({\mathbb{T}}^{2},{\mathbb{Z}})$ vanishes by [27, (1.14)], whence triviality by [57, Proposition, p.2] again.
@@
    $C({\mathbb{S}}^{3}_{\theta})$ is finite over its center
@@
[…]

### v4

@@ 3 Local-Triviality Dimension and Fields
and satisfy the conditions that each $\alpha_{i,s}^{k}$ is a projection and $\sum\limits_{s=1}^{\operatorname{\mathrm{qt}}}\alpha_{i,s}^{*}\alpha_{i,s}$ is invertible on each $U_{i}$. As before, we have that the $\alpha_{i,s}$ are in the generating spectral subspace of the ${\mathbb{Z}}/k$-action by design. However, each $\alpha_{i,s}$ is only a section on $U_{i}$, not the full space $X$.
@@
A partition of unity argument (see [35, Theorem 36.1]) allows us to rescale the operators $\alpha_{i,s}\in\Gamma({\mathcal{M}}|_{U_{i}})$ into global sections $\phi_{i}^{1/2}\alpha_{i,s}\in\Gamma(M)$, where the $\phi_{i}$ are nonnegative scalar-valued continuous functions on $X$, supported in $U_{i}$, with $\phi_{i}|_{V_{i}}\equiv 1$ and $\sum\limits_{i=0}^{\operatorname{ix}}\phi_{i}\equiv 1$. Note that each $\phi_{i}$ is a central element of $\Gamma({\mathcal{M}})$ and is fixed by the action of ${\mathbb{Z}}/k$, so multiplying by $\phi_{i}^{1/2}$ has not removed any of the key properties of $\alpha_{i,s}$. Because $1\leq s\leq\operatorname{\mathrm{qt}}$ and $0\leq i\leq\operatorname{ix}$, we have produced a list of $\operatorname{\mathrm{qt}}\cdot(\operatorname{ix}+1)$ normal operators $\alpha_{i,s}$ in the generating spectral subspace of the ${\mathbb{Z}}/k$-action such that $\sum\limits_{i,s}\alpha_{i,s}^{*}\alpha_{i,s}$ is (globally) invertible. By Proposition 3.5, we have that
@@
The implication a $\Longrightarrow$ b follows from the upper bound 3-4 on $\mathrm{dim}^{{\mathbb{Z}}/k}_{\mathrm{WLT}}(\Gamma({\mathcal{M}}))$, where we note that $\operatorname{ix}({\mathcal{M}},\triangleright)$ is always finite, and the full spectrum condition of a guarantees that the definition of $\operatorname{\mathrm{qt}}$ does not involve dividing by zero. Moreover, b $\Longrightarrow$ c follows from the general result [20, Theorem 3.8]. As for c $\Longrightarrow$ a, freeness means by [34, Theorem 8.1.7] that the ${\mathbb{Z}}/k$-grading on $A:=\Gamma({\mathcal{M}})$ induced by the action is strong:
@@
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A)=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A)=1\quad\text{and}\quad\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{SLT}}(A)=\infty.$
@@
Example 3.7 already argues that none of the dimensions can vanish, and the upper bound of $1$ for the ordinary (hence also weak) LT dimension follows from Theorem 3.15 item a by the selfsame example: $\operatorname{ix}({\mathcal{M}},\triangleright)=1$, since a cover by two punctures of ${\mathbb{P}}^{1}$ trivializes everything in sight.
As for the infinitude of the strong LT dimension, recall first that the action at $x\in{\mathbb{P}}^{1}$ is conjugation by the self-adjoint (hence involutive) unitary having $x$ (regarded as a line in ${\mathbb{C}}^{2}$) and the orthogonal complement $x^{\perp}\leq{\mathbb{C}}^{2}$ as its $1$- and (1)(-1)-eigenspaces, respectively.
@@
${\mathbb{P}}^{1}\cong{\mathbb{S}}^{2}\quad\ni\quad x\xmapsto{\quad}\big(\text{pair of $a_{i}(x)$-eigenspaces}\big)$ $\in\quad{\mathbb{S}}^{2}/(\text{antipodal ${\mathbb{Z}}/2$})$ (3-5)
$\cong\quad\text{real projective plane }{\mathbb{R}}{\mathbb{P}}^{2}.$
Because (by oddness) each eigenspace of $a_{i}$ is orthogonal to each eigenspace of $u$, 3-5 coincides with the canonical surjection ${\mathbb{S}}^{2}\to{\mathbb{R}}{\mathbb{P}}^{2}$ nowhere. A continuous lift ${\mathbb{S}}^{2}\to{\mathbb{S}}^{2}$ thereof (which does exist by [35, Lemma 79.1], as ${\mathbb{S}}^{2}$ is simply-connected) will thus have no coincidences with either the identity or the antipodal map, once more violating [59, Corollary 1.24]. $\blacksquare$
@@
by projectivity, as in the proof of [12, Proposition 5.4], via [46, Proposition 2.10]. Define $a:=\sum\limits_{i=0}^{n}\overline{\rho}_{i}(t\otimes 1)$. Note that $\pi_{x}(a)$ is invertible, so we may choose an element $b\in A$ such that $\pi_{x}(b)=(\pi_{x}(a))^{-1}$. That is, both $\|ab-1\|_{x}=0$ and $\|ba-1\|_{x}=0$. By [48, Proposition 1.2], for each $c\in A$ the mapping
@@
Consider the antipodal ${\mathbb{Z}}/2$-action on the $\theta$-deformed spheres $C(\mathbb{S}^{3}_{\theta})$, which form a continuous field. The Borsuk-Ulam theorem implies that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(C({\mathbb{S}}^{3}))=3$. On the other hand, for $\theta\not\in{\mathbb{Z}}$ the proof of [40, Theorem 2.10] shows that if $C({\mathbb{S}}^{3}_{\theta})$ is noncommutative, with $z_{j}=x_{j}+iy_{j}$, then
@@
We conclude that the weak local-triviality dimension function is upper semicontinuous, but not continuous, in the parameter $\theta$. $\blacklozenge$
@@
where ${\mathbb{S}}^{1}$ is a maximal torus: see [30, Theorems 1 and 2] and [29, Theorem III.1.11.1], for $G=SU(2)$, as recalled in [52, Introduction].
@@
    $SU(2)\xrightarrow{\quad}PSU(2)\cong SO(3);$ (3-6)
@@
    resulting from realizing ${\mathbb{C}}^{n}$ as the irreducible $n$-dimensional representation of $SU(2)$ (see [55, Theorem VIII.4.2]).
@@
On the other hand, $\sigma$ acts on ${\mathbb{C}}^{n}$ with eigenvalues $\pm\lambda$, with $\lambda=1$ or $\lambda=i$ depending on the parity of $n$, and with the $(\pm\lambda)$-eigenspaces of dimension as close to equal as possible. It follows that $\sigma$ induces the usual gauge $({\mathbb{Z}}/2)$-action on $M_{n}$ resulting from the latter’s realization as the graph algebra of a length-$n$ directed path. Finally, [12, Proposition 4.5] shows that
@@
The $A_{\theta}$ are the fibers of a continuous field by [51, Theorem 8.13 and Example 10.2], and all admit ergodic actions (see [21, p.537]) by the ordinary 2-torus ${\mathbb{T}}^{2}\cong({\mathbb{S}}^{1})^{2}$. This constitutes a continuous field of actions in the sense of [48, Definition 3.1]. Note the following.
@@
  • On the other hand, for non-integral values of $\theta$, none of the dimensions can vanish. The ${\mathbb{T}}^{2}$-action is ergodic, so the eigenspaces for the characters of ${\mathbb{T}}^{2}$ are all 1-dimensional (see [60, Theorem 1]). Therefore an equivariant morphism
@@
    $C({\mathbb{G}})\xrightarrow[]{\quad}\mathrel{\mkern-14.0mu}\rightarrow A\xrightarrow[]{\quad}\mathrel{\mkern-14.0mu}\rightarrow C_{r}({\mathbb{G}}),$
@@
with full equivalence when $\mathbb{G}$} is of Kac type [37, post Proposition 1.7.9].
@@
Recall from [61, Corollary 4.1] that $\varphi$ uniquely retrieves a completely positive contractive order-zero [61, Definition 2.3] map
@@
Note that $\widetilde{\varphi}$ is $\mathbb{G}$}-equivariant in the usual sense, and by [61, Theorem 3.3] it is a scaling of a morphism $C({\mathbb{G}})\to A$ by the $\mathbb{G}$}-invariant, positive, non-zero element $h:=\widetilde{\varphi}(1)\in A$. Since ergodicity of the action implies that $h$ is a scalar, we can divide by it to recover a $\mathbb{G}$}-equivariant $C^{*}$-morphism $C({\mathbb{G}})\to A$.
Observe also that a $\mathbb{G}$}-equivariant $C^{*}$-morphism is by necessity one-to-one on the dense Hopf *-algebra [37, §1.6] ${\mathcal{O}}({\mathbb{G}})\leq C({\mathbb{G}})$: the $\mathbb{G}$}-invariant state
@@
restricts to $h_{{\mathbb{G}}}$ along the assumed morphism $C({\mathbb{G}})\to A$, and $h_{{\mathbb{G}}}$ is faithful on ${\mathcal{O}}({\mathbb{G}})$ [37, Corollary 1.7.5]. If the action is 0-dimensional, then, the multiplicity of a $\mathbb{G}$}-representation $\rho$ in $A$ is precisely $\dim\rho$. For ergodic actions of Kac groups this is also an upper bound for that multiplicity [7, Theorem 17] (see also [60, Theorem 1] for classical compact groups), hence the implication a $\Longrightarrow$ d for Kac-type $\mathbb{G}$}. $\blacksquare$
@@
of ergodic ${\mathbb{T}}^{n}$-actions in Example 3.20 defined over the skew-symmetric matrices $\theta$ has local-triviality dimension equal to 0 precisely when $\theta$ is integral. Moreover, it has local-triviality dimension $\infty$ elsewhere. This proves a stronger claim than used in Example 3.20.
@@
  • if the weak local-triviality dimension (or rather, any one of the local-triviality dimensions) of an action is finite, then [20, Theorem 3.8] shows that the action is free;
@@
respectively for some primitive $q^{th}$ root of unity $\zeta$ (see [21, proof of Proposition 12.2, (12.7)]). This in fact realizes $M_{q}$ as a cocycle twist [17, §2.2] $C^{*}(\Gamma,\sigma)$ of the group algebra of
@@
by [53, Theorem 9.74 and Theorem following Example 9.65]. Next, [34, Theorem 8.1.7] once more shows that the induced grading by the Pontryagin-dual group $\Gamma$ is strong:
@@ 4 Rational tori and spheres
However, if some particular $t_{j}$ equals 0, then $z_{j}=t_{j}u_{j}=0$ is never changing, so the function $g$ must not depend on $u_{j}$. That is, $g$ is in the $C^{*}$-algebra generated by the other coordinate functions. This perspective extends to the $\theta$-spheres and quantum tori, as in [36, Theorem 2.5]. We recall this result and introduce some notation below.
@@
    for the “first-quadrant” portion of the (n1)(n-1)-sphere.
@@
    Each ${\mathbb{S}}^{n-1}_{++,F}$ is the interior of an (|F|1)(|F|-1)-dimensional simplex, and the closed simplex ${\mathbb{S}}^{n-1}_{+}$ is the disjoint union of the various open simplices ${\mathbb{S}}^{n-1}_{++,F}$.
@@
    ${\mathbb{S}}^{n-1}_{+,F}$ $=\,\,\overline{{\mathbb{S}}^{n-1}_{++,F}}$
    $=\,\,\left\{(t_{i})\in{\mathbb{S}}^{n-1}_{+}\ |\,\,\forall i\not\in F,\ t_{i}=0\,\right\}.$
For a skew-symmetric matrix $\theta\in M_{n}({\mathbb{R}})$, recall from [36, Theorem 2.5] that
@@
For k=2n1k=2n-1, each generator $z_{j}=x_{j}+iy_{j}$ of $C({\mathbb{S}}^{2n-1}_{\theta})$ is normal and odd, so the identity $1=\sum\limits_{j=1}^{n}z_{j}z_{j}^{*}=\sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}$ shows that the weak and plain local-triviality dimensions are at most 2n12n-1. A similar argument applies to k=2n2k=2n-2, with the final generator $z_{n}=x_{n}$ self-adjoint. $\blacksquare$
@@
Every $\theta$-sphere has $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))\geq k$. If each entry of $\theta$ is a rational number with odd denominator, then in fact $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))=k$.
@@
The noncommutative Borsuk-Ulam theorem [40, Corollary 3.8] implies that if $m<k$, there is no $\mathbb{Z}/2$-equivariant morphism $C({\mathbb{S}}^{m}_{\phi})\to C({\mathbb{S}}^{k}_{\theta})$. This applies in particular to the case when the domain sphere is commutative. Nonexistence of an equivariant morphism $C({\mathbb{S}}^{m})\to C({\mathbb{S}}^{k}_{\theta})$ for $m<k$ is equivalent to the claim $\mathrm{dim}^{\mathbb{Z}/2}_{\mathrm{SLT}}(C({\mathbb{S}}^{k}_{\theta}))\geq k$.
Now, assume that each entry of $\theta$ is a rational number with odd denominator, and suppose k=2n1k=2n-1 is odd. There exists an odd positive integer $M$ such that in the associated quantum torus $C(\mathbb{T}^{n}_{\theta})$, the elements $u_{1}^{M},\ldots,u_{n}^{M}$ are central. The functions $f_{i}:{\mathbb{S}}^{n-1}_{+}\to C(\mathbb{T}^{n}_{\theta})$ defined by
@@
written in lowest terms. From [47, Theorem 3.12] (or [21, Proposition 12.2]) we see that, in this case, there is a rank-$q$ vector bundle $\mathcal{E}$} on ${\mathbb{T}}^{2}$ such that
@@
This is also recalled in [27, §2.5], which is perhaps more accessible. The center
@@
is generated by the two unitaries $u^{q}$ and $v^{q}$ [21, Corollary 12.3], and $A_{\theta}$ is a bundle of $q\times q$ matrix algebras over that center. This also makes $A_{\theta}$ into a rank-$q^{2}$ Azumaya algebra (see [33, §§13.7.6 and 13.7.13], [9, §§III.1.3 and III.1.4], [1, Definition 5.4.17], etc.): it is finitely-generated projective over its center $Z$, with
@@
    The discussion following [27, Exercise 2.7] claims that the rank-$q$ vector bundle $\mathcal{E}$} on ${\mathbb{T}}^{2}$ yielding 4-3 is flat, which it cannot be. A flat bundle (see [28, §1.2]) admits a connection with vanishing curvature, or local trivializations with locally constant transition functions, and flat bundles have vanishing Chern classes by [28, Proposition II.3.1]. In contrast, the Chern class
@@
    is identified in [47, Proposition 3.8 and Theorem 3.9] with an integer $a$ for which $ap+bq=1$ (the twist of [47, p.299] is a-a). In particular, that Chern class cannot vanish. Note that the vector bundle $\mathcal{E}$} is only unique up to the tensor product with a line bundle [15, Théorème 9], so $a$ is only unique modulo $q$. This is also compatible with the fact that $A_{\theta}$ only depends on the coset of $\theta$ modulo $\mathbb{Z}$}.
    Additionally, vector bundles over compact connected Riemann surfaces are completely determined by their ranks and degrees (i.e. Chern classes $c_{1}$) by [58, Proposition, p.2]. Flat bundles on ${\mathbb{T}}^{2}$, for that reason, are topologically trivial. They need not be holomorphically trivial, but ‘topologically’ is what matters in the present context.
@@
    What does hold, and what was perhaps meant by the flatness claim in the paragraph preceding [27, Exercise 2.8], is that the endomorphism bundle $\mathcal{E}$$\otimes{\mathcal{E}}^{*}$ (recovering $A_{\theta}$ as the section algebra 4-3) is flat, and in fact is topologically trivial as a vector bundle (but not as a bundle of algebras!). The first Chern class $c_{1}({\mathcal{E}}\otimes{\mathcal{E}}^{*})\in H^{2}({\mathbb{T}}^{2},{\mathbb{Z}})$ vanishes by [28, (1.14)], whence triviality by [58, Proposition, p.2] again.
@@
    The center of $C({\mathbb{S}}^{3}_{\theta})$ is
@@
    $=\left\{f:[0,1]\xrightarrow{\emph{cont.}}C^{*}(u^{q},v^{q})\cong C({\mathbb{T}}^{2})\quad|\quad f(0)\in C^{*}(u^{q}),\ f(1)\in C^{*}(v^{q})\right\}$